OpenStudy (owlet):

Alizarin yellow R, \(Ka=7.9×10^{12}\), is yellow in its protonated form (HX) and red in its ionized form \((X^−)\). At what pH will alizarin yellow R be a perfect orange color? I have no idea how to do this. I tried finding the pH of alizarin yellow, which is 11.1. Then what's next?

OpenStudy (owlet):

I think this one helps:

OpenStudy (owlet):

@Rushwr

OpenStudy (rushwr):

According to the indicators' colour change the pH at perfect orange is around 4 right?

OpenStudy (mortonsalt):

Isn't the "perfect orange" color the moment where Alizarin shifts from its protonated form to its ionized form?

OpenStudy (mortonsalt):

\[HX \rightarrow H^+ X^-\] \[K_a = \frac{[H^+][X^-]}{[X^-]}\] \[K_a = [H^+]\] Then solve for pH.

OpenStudy (mortonsalt):

I'm not entirely sure, though, haha! Don't take my word for it.

OpenStudy (rushwr):

That's not right ! \[K _{a} = \frac{ [H ^{+}] [X ^{-}] }{ [HX] }\]

OpenStudy (rushwr):

So basically this is a weak acid

OpenStudy (rushwr):

@owlet Are you given with the concentration ?

OpenStudy (mortonsalt):

Hang on, \[[HX] = [X^-]\] because it's at equilibrium at the perfect orange point.

OpenStudy (rushwr):

yeah that's right so Ka = H^+

OpenStudy (rushwr):

pH = -log {H^+]

OpenStudy (mortonsalt):

So where did I go wrong? o:

OpenStudy (owlet):

sorry late response. No, I'm not given any concentrations :

OpenStudy (rushwr):

@mortonsalt Ur Ka expression is wrong check it again

OpenStudy (mortonsalt):

You don't need it, really. \[pH = pK_a = -\log(K_a)\]

OpenStudy (rushwr):

Yeah since u gave it only I said it's not right

OpenStudy (mortonsalt):

That's the expression for when I've already substituted the value in.

OpenStudy (mortonsalt):

But yeah, anyway! :) Hope this helps.

OpenStudy (owlet):

I already did that, pH = -log(7.9x10^-12) = 11.1 , when it is on the yellow interval but how about when it is already on orange?

OpenStudy (owlet):

should I subtract the pH from it from the perfect orange color?

OpenStudy (mortonsalt):

No, that's not necessary. Look at it this way: Since [H+] = Ka Then pH = pKa

OpenStudy (mortonsalt):

This is at the equilibrium point, like how we got it previously.

OpenStudy (mortonsalt):

STEP 1 \[K_a = \frac{[H^+][X^-]}{[HX]}\] STEP 2 Since we know that we're looking for the "perfect orange", this is the point where the two "colors" are at equilibrium. Therefore we can say that \[[HX] = [X^-]\] We substitute the first equation to the second equation and we get \[K_a = \frac{[H^+][X^-]}{[X^-]}\] Are you okay so far?

OpenStudy (rushwr):

@owlet do u have any way to check the answer?

OpenStudy (mortonsalt):

STEP 3 This means that \[K_a = [H^+]\] From the equation above, we can say that \[pH = pK_a\]

OpenStudy (owlet):

yeah, i guess. this may sound dumb, but how do you know that the perfect orange is where the colors are at equilibrium? yeah. it is right 11.1 :) I just wanted to understand this topic.

OpenStudy (mortonsalt):

Red + yellow?

OpenStudy (mortonsalt):

And it says so on the picture too that right at the middle when [HX] is equal with [X^-].

OpenStudy (owlet):

oh, I missed the red color ! that's why, okay thank you for the both of you :D

OpenStudy (rushwr):

@owlet Here there's nothing to do . As we can see in the diagram u attached we can see the perfect orange is obtained when HX is approximately equal to X^- So that's how u get it .

OpenStudy (mortonsalt):

Just think of it this way too: It's a reversible process. The protonated ones turn into ionized ones and vice versa. At ONE POINT, there is an equal number of protonated ones to ionized ones.

OpenStudy (mortonsalt):

This is the equilibrium point.

OpenStudy (rushwr):

 Therefore we cancel off HX to X^-